Factor the expressions.
x^2+x-12
x^2+11x+30
x^2+3x+2

Answer by ewatrrr(24785)

 
Hi,
x^2+x-12 = (x+4)(x-3)
x^2+11x+30 = (x+5)(x+6)
x^2+3x+2 = (x+1)(x+2)
check the correctness of the factoring by multiplying them using the FOIL method
F First terms
O Outside terms
I Inside terms
L Last terms